# Cyber-Twin: Digital Twin-boosted Autonomous Attack Detection for Vehicular Ad-Hoc Networks

### 09 Feb 2024 v3

@@ V Conclusion
In conclusion, our research has successfully introduced a cyber-twin framework that significantly enhances the security of Road Side Units (RSUs) in Vehicular Ad-hoc NETworks (VANETs). We have developed a robust method for real-time vehicular network monitoring and management by integrating digital twin technology with advanced AI models. This framework adeptly addresses VANETs’ complexities, notably in vehicle-to-infrastructure communications. Our findings demonstrate substantial resource utilization and attack detection improvements, outperforming existing solutions. The cyber-twin framework’s ability to create accurate digital replicas of RSUs has been crucial in effectively detecting and mitigating cyber threats, reinforcing communications’ integrity and reliability. Furthermore, the framework significantly advances green communication practices by enhancing RSU computational efficiency, contributing to energy efficiency and hardware longevity. This aligns with the shift towards sustainable, eco-friendly technologies in intelligent transportation systems and smart cities. Overall, our cyber-twin framework sets a new benchmark in developing secure, sustainable, and intelligent transportation systems, addressing key security challenges while paving the way for a future of safer and greener smart cities.

### 15 Mar 2024 v4

@@ V Conclusion
Our research has successfully introduced a novel framework that significantly enhances the security of RSUs in VANETs. We have developed a robust real-time vehicular network framework that integrates cyber twins with advanced AI models. The framework adeptly addresses the complexities of VANETs, particularly in vehicle-to-infrastructure communications. Our research findings demonstrate substantial resource utilization and attack detection improvements, which outperform existing solutions. The framework’s cyber twins of RSUs help detect and mitigate cyber threats, ensuring communication integrity and reliability. Our framework also contributes to green communications by improving the computational efficiency of RSUs, reducing the average delay of data transmission, and leading to increased energy efficiency and extended hardware durability. It has also achieved an optimal balance between resource consumption and high attack detection effectiveness, with a defined twinning rate range of seventy-six to ninety per cent. These advancements underscore our commitment to developing sustainable, secure, and resilient vehicular communication systems for the future of smart cities. Overall, our framework sets a new benchmark in developing secure and green ITS, addressing security challenges while paving the way for a future of safer and greener smart cities.